OpenStudy (anonymous):

How would the expression x^2-4 be rewritten using Difference of Squares? A. (x-4)^2 B.(x+2)(x-2) C.(x+2)^2 D.(x+4)(x-4)

rishavraj (rishavraj):

(a^2 - b^2) = (a + b)(a - b)

OpenStudy (anonymous):

what...lol

rishavraj (rishavraj):

\[(a^2 - b^2) = (a + b)(a - b)\] so \[(x^2 - 4) = (x^2 - 2^2)\]

OpenStudy (anonymous):

B

rishavraj (rishavraj):

yeah :)

OpenStudy (anonymous):

thank you :p :)